OpenStudy (anonymous):

help!! which equation represents the line passing through the points (3,2) and (-9,6)

OpenStudy (e.mccormick):

Have you tried finding the slope then using point slope to find the line?

OpenStudy (anonymous):

yea i got 3x+y=9

OpenStudy (ryan123345):

First find the slope.

OpenStudy (anonymous):

how do u do tht ?

OpenStudy (e.mccormick):

How did you get 3x? Look at your slope formula:\[m=\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (e.mccormick):

AKA rise over run or change in y over change in x.

OpenStudy (anonymous):

oh so should the 9 be -?

OpenStudy (e.mccormick):

Well, whatever it is, you subtract it. So if it is negative, it gets added. \(a - (-b)=a+b\)

OpenStudy (e.mccormick):

Here is another interesting fact. It does not matter which you choose as \((x_1,y_1)\) and \((x_2,y_2)\). For example: \[\frac{2-6}{3-(-9)}~\mathrm{and}~\frac{6-2}{(-9)-3}\]both give the same answer. Once you solve the slope properly, put it into a point-slope formula. Then you can get the answer.